A grandma's age is 9 years less than 7 times her grand daughter's age. The sum of their ages is 95. Find the age of each.
Answer by aquacean(7)

Orthodox method:
a = grand daughter's age
95 = a + 7a - 9
104 = 8a
104 / 8 = a
a = 13
b = grandmother's age
b = (7)(13) - 9
b = 82

My method:
I want to explain but I feel sleepy now
1.) 95 + 9 = 104
2.) 104 / 8 = 13
3.) 13 = grand daughter's age
4.) 13 * 7 = 91
5.) 91 -9 = 82 = grandma's age